OpenStudy (anonymous):

2x^2 -3x=x^2+10 solve by factoring

OpenStudy (skullpatrol):

Any ideas? :-)

OpenStudy (anonymous):

2x^2-3x=x^2+10 2x^-x^2=10+3x x^2=10+3x x=5

OpenStudy (anonymous):

i am right?

OpenStudy (anonymous):

i think there should be two answers for x..

OpenStudy (danjs):

i think they want you to move it into standard form first y = ax^2 + bx + c = 0

OpenStudy (danjs):

x^2 - 3x - 10 = 0 (x )*(x ) = 0

OpenStudy (anonymous):

ok guys) i will try to do for this formula)

OpenStudy (anonymous):

(x-5)(x+2)

OpenStudy (danjs):

use factors of 10 to try and create a -3 in the middle

OpenStudy (danjs):

+2 and -3

OpenStudy (danjs):

put those in and expand to see if it worked.. (x + 2)*(x - 3) = 0

OpenStudy (anonymous):

oh okay thanks guys!!! :)

OpenStudy (danjs):

welcome, so either of those ( ) can be zero to make it true, x=-2 or x = 3

OpenStudy (danjs):

or you can always apply the quadratic formula and see what th evalues are to give a hint as to the factors

OpenStudy (anonymous):

thank you very much